OpenStudy (ilovecomputers1):

When looking at a rational function, Charles and Bobby have two different thoughts. Charles says that the function is defined at x = −2, x = 3, and x = 5. Bobby says that the function is undefined at those x values. Describe a situation where Charles is correct, and describe a situation where Bobby is correct. Is it possible for a situation to exist where they are both correct? Justify your reasoning.

OpenStudy (ilovecomputers1):

Y=X is an example of a case where Charles would be correct i think. But i dont get bobby

OpenStudy (anonymous):

Any rational function that has a factor of \((x-k)\) in the denominator will be undefined for \(x=k\). For examples, \(\dfrac{1}{x+2}\) is undefined when \(x=-2\). So any function that looks of this form will work for Charles' claim. The opposite is true for Bobby. You can use any function that doesn't have such a factor in the denominator. One thing to note for Charles is that a function like \[\frac{x^2-4}{x+2}\] will not work, even though you can remove the discontinuity at \(x=-2\) by dividing: \[\frac{x^2-4}{x+2}=\frac{(x-2)(x+2)}{x+2}=x-2\] HOWEVER, this algebra only works for \(x\not=-2\), so in fact you would have \[\frac{x^2-4}{x+2}=\begin{cases}x-2&\text{for }x\not=-2\\\text{undefined}&\text{for }x=-2\end{cases}\]